Recent Log Entries
iocsoftnl10rf restoration failed
iocsoftnl10rf restored
iocnl11aclrm restored
New injector record?
Injector is pretty warm. Tunnel work says it feels like 90F, like it agrees with the thermocouple.
Tunnel worker says it feels like 90F in the injector, agrees with thermocouple.
Steel is not a great thermal conductor, time would tell but wouldn't be surprised to find small increase in vacuum in the baked beam line (fig. 2) as temperature increases from ~70 F to ~90 F in tunnel.
Can we turn the AHU's back on? Is there a method to monitor them remotely?
iocipmk203 restored
Operator In Training SIGN IN
Operator SIGN IN
Operator SIGN IN
HYST AREA 6-2-4 Cycles Magnets
HYST AREA Version 6-2-4 has FAILED to cycle:
MBC1H01V -- BS04B10 -- sb4mag -- Failed to set after 2 cycle attempts
MBDAR08V -- NL01B03 -- nl6mag -- Failed to set after 2 cycle attempts
HYST AREA Version 6-2-4 has successfully cycled the following:
Area.Injector.PGun
Area.Injector.Linac
Area.Injector.Chicane
Area.Linacs.North
Area.Linacs.South
Area.Arcs.Arc1.Spreader
Area.Arcs.Arc1.Arc
Area.Arcs.Arc1.Recombiner
Area.Arcs.Arc1.Extraction
Area.Arcs.Arc2.Spreader
Area.Arcs.Arc2.Arc
Controls recovery after NL breaker work
In support of Sasa and Robert from cryo the following iocs and associated CAMAC crates were powered down and back up: iocnl1b, iocnl2, iocnl4. All control hardware was successfully restored. Thanks to S. Higgins for software support.
Box Supplies Were Cycled
The following box supplies were CYCLED by version 2-7 of the Box Supply Cycling Script:
ARC1A Ramp Rate = 8.203 A/s
ARC2A Ramp Rate = 6.851 A/s
ARC3A Ramp Rate = 9.974 A/s
ARC4A Ramp Rate = 10.852 A/s
ARC5A Ramp Rate = 10.298 A/s
ARC6A Ramp Rate = 9.271 A/s
ARC7A Ramp Rate = 10.627 A/s
ARC8A Ramp Rate = 12.126 A/s
ARC9A Ramp Rate = 9.254 A/s
ARCAA Ramp Rate = 12.088 A/s
LAM1C Ramp Rate = 11.764 A/s
LAM3C Ramp Rate = 11.796 A/s
MYA2T01 Ramp Rate N/A
MYA4T01 Ramp Rate N/A
MYA6T01 Ramp Rate N/A
MYA8T01 Ramp Rate N/A
MYAAT01 Ramp Rate N/A
RSEP8A Ramp Rate = 10.630 A/s
RSEP9A
Follow-up Re: Follow-up Re: NLinac heat
Gun HV Turned Off
Gun HV Turned Off
FSD Masking Tool: Successful Destination masking to: BSY
iocnl4 restored
iocnl2 restored
COMMENT ON MPS add FSD Master output signal to NL and SL for RF systems use
SL fiber is now patched in from FSD Master to SL22B14. Fiber coil is on top of the rack and ready for use (plugged in at FSD master end). Orange fiber is for use, blue fiber is spare. NL patched through to NL zone 1 Junction box. still needs to be routed to zone 22. Note pic attached for reference.